OpenStudy (anonymous):

Let G= {a+b*2^(1/2) | a,b Element of Q} Q is the set of rational numbers. Prove that G is a group under addition.

myininaya (myininaya):

ok we need to show G is closed, associative, has identity, and has inverse. *CLOSURE! so we need to show when we take to elements of set G then we still get an element in set G. so Let a,b,c,d be rational. so we have that a+b*2^(1/2) is element of G and c+d*2^(1/2) is element of G so when we add: (a+b*2^(1/2))+(c+d*2^(1/2)) =(a+c)+(b+d)*2(1/2) so we need a+c to be rational and b+d to be rational if you take a rational number and add it to another rational number you still have a rational number so G is closed! now step 2 show that G is associative

OpenStudy (anonymous):

Ok, I think I am starting to get it.

myininaya (myininaya):

*Associative let a,b,c,d,e,f be rational. so we have that a+b*2^(1/2) is element of G c+d*2^(1/2) is element of G e+f*2^(1/2) is element of G so we need to show [(a+b*2^(1/2))+(c+d*2^(1/2))]+(e+f*2^(1/2))= (a+b*2^(1/2))+[(c+d*2^(1/2))+(e+f*2^(1/2))] this step is not hard to do

myininaya (myininaya):

we have the top part = (a+c)+(b+d)*2^(1/2)+(e+f*2^(1/2)) =(a+c+e)+(b+d+e)*2^(1/2) bottom part= (a+b*2^(1/2))+(c+e)+(d+f)*2^(1/2) =(a+c+e)+(e+d+f)*2^(1/2) since addition is commutative and associative then G is associative

myininaya (myininaya):

*Identity let a,b rational then a+b*2^(1/2) is element of G so we have a+b*2^(1/2)+?=a+b*2^(1/2) if ? is an element of G, then G is the identity property

OpenStudy (anonymous):

Thank you so much for this.

myininaya (myininaya):

if ?=0, then the equation holds but is 0 and element of G 0 and 0 are rational numbers so 0+0*2^(1/2) is element of G but 0+0*2^(1/2)=0 so this must mean 0 is element of G and therefore has the identity property

myininaya (myininaya):

last step is Inverse

myininaya (myininaya):

*Inverse again let a,b rational then a+b*2^(1/2) is element of G now we want to find some element in G,?, such that (a+b*2^(1/2))+?=0 if we can ? is in G, then we have the inverse property additive inverse of a is -a additive inverse of b*2^(1/2) is -b*2^(1/2) -a and -b are both rational so therefore -a-b*2^(1/2) is element of G and therefore has inverse property

myininaya (myininaya):

*in conclusion since we have all four of the above, we have that G is a group under addition.

myininaya (myininaya):

was that fine? :)

OpenStudy (anonymous):

oh yeah. thank you.

myininaya (myininaya):

np :)

OpenStudy (zarkon):

you could have also use the subgroup test to show that G is a subgroup of the group \[(\mathbb{R},+)\] subgroups are groups